Littlefeet Companies is currently trading at $43.75 a share. The company has 18,000,000 shares outstanding and earned $36,000,000 last year. It is in the Consumer Retail (Leisurewear) sector - companies in that sector generally trade at a 15.5 P/E multiple. You are in investor that generally invests in companies that are trading at P/E multiples lower then those of their overall sector. Would be likely to buy Littlefeet Companies? Choose the answer with the most accurate description of your answer and the reasoning behind the answer.

a.Yes because Littlefoot's share price is higher then the sector multiple.

b.No, the P/E of Littlefeet is higher then the sector multiple.

c.Yes because Littlefeet Companies has a higher P/E then its sector

d.No because Littlefoot's P/E is lower then the sector multiple
Answer #1

b.No, the P/E of Littlefeet is higher then the sector multiple.

because littlefeet P/E=43.75/(36000000/18000000)=21.9 which means higher than sector's 15.5

the above is answer..